OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (jojo4eva):

the slope would be -14

OpenStudy (jojo4eva):

-6-8/-3-(-4) = -14/-3+4=-14/1=-14

OpenStudy (anonymous):

Oooh Thank you so much!!

OpenStudy (anonymous):

It is - 14

OpenStudy (jojo4eva):

no problem if you need any help just tag me